OpenStudy (anonymous):

An angle measures 12 degrees less than three times its supplement. Find the measure of the angle

OpenStudy (anonymous):

its an equarion

OpenStudy (ddcamp):

Supplementary angles add up to 180° Let's start with having two angles a and b. a is the angle you're trying to find, and b is the supplementary angle: a + b =180° a = 3b - 12 Can you solve it from here?

OpenStudy (anonymous):

will it be equal to each other

OpenStudy (ddcamp):

No, the two equations aren't equal to each other right now. You could subtract b from both sides in the first equation to give: a = 180 - b Then set the equations equal to each other: 3b-12 = 180-b

OpenStudy (anonymous):

i am confused

OpenStudy (ddcamp):

Do you know how to solve systems of equations?

OpenStudy (anonymous):

yess

OpenStudy (ddcamp):

The two equations are: a + b =180° a = 3b - 12 How would you start to solve that?

OpenStudy (anonymous):

so you will cancel the a

OpenStudy (anonymous):

will it be like 2b=168

OpenStudy (anonymous):

wont the answer be 84??

OpenStudy (ddcamp):

I would start to solve the system by replacing the a in the first equation with 3b-12 (substitution), giving: 3b - 12 + b = 180 4b - 12 = 180 4b = 192 b = 48